40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Feb 13 16:50:17 2011, kiwamu, Update, ASC, IP_ANG was at wrong ADC xmen.jpg
    Reply  Mon Feb 14 12:37:23 2011, kiwamu, Update, ASC, no signal from IP_ANG_Seg1 
       Reply  Tue Feb 15 06:15:07 2011, Suresh, Update, ASC, no signal from IP_ANG_Seg1 Y-end_beam.JPG
Message ID: 4296     Entry time: Tue Feb 15 06:15:07 2011     In reply to: 4287
Author: Suresh 
Type: Update 
Category: ASC 
Subject: no signal from IP_ANG_Seg1 

[Valery, kiwamu, Jenne, Suresh]

    I first interchanged the two QPD's on the Y end table to see if the problem QPD related.  Exchanging the units did not make any difference.  The problem therefore had to be in the cables or the circuit boards in 1X4

    We traced the signals pertaining to the IP_ANG QPD ( "Initial Pointing Beam") using  Jay's wiring diagram (pages 2 and 5 of 7).  We noted that while the signals were available on all Segments till the Monitors (Lemo) on 1X4-2-2A card, two of the lines did not reach the output of the cross connect 1X4-B8.  We checked card to make sure that the signals were indeed reaching the back plane of the 1X4-2 chassis using a D990612 extension board.  The card was found to be okay.  We therefore suspected that the cable (CAB_1X4_?) going from the card to the cross connect 1X4-B8 was faulty.  Indeed visual inspection showed that the crimping of the connector was poor and weight of the cable had put further strain on the crimping.  

   I changed the 64-pin connector on the 1X3-2-2A side of the cable. 

When I connected everything back together the problems persisted. Namely the lines P1-1A  (Segment 1 high) and P1-2C (Segment 2 Low) were floating They were not reaching points 2T and 3T respectively on the output of the cross connect.

   I therefore replaced 1X4-B8 with a similar unit which I found in one of the shelves along the East (Y) arm. 

I then checked with the StripTool to make sure that all the quadrants are showing similar response to a flashlight on the QPD.   All Segments are working fine now. Currently the IR Initial Pointing beam reaches the QPD but is not centered on it. 

I did not attempt to center it since the beam appeared to be clipped and may anyway require repositioning.

JD: We need to meditate on where this beam could be getting clipped.  Suresh and I checked that it's not on the viewport on the beam's way out of the ETMY chamber by seeing that the beam is far away from the edges of the viewport, and also far away from the edges of the black beamtube between the viewport and the table.  Suresh mentioned that the clipping nature of the IP_ANG beam sometimes goes away.  I don't know if this is the same clipping that Kiwamu might be seeing with the main beam, or if this is separate clipping just with the IP beam, after it's been picked off.  I suspect it's the same as what Kiwamu is seeing....maybe when we move PZT1, we clip on one of the MMT mirrors or PZT2??  If this is true, it's a total pain since we might have to vent if we can't steer around it.

Y-end_beam.JPG

 

ELOG V3.1.3-